# Waring and Waring–Goldbach subbases with prescribed representation function

### v1

@@ 5.2.2. Integral over minor arcs
For $\ell\geq 2^{k}$, we have
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{c(k)+2^{k}}.$
@@
The integral $\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha$ counts the number of solutions to the Diophantine equation (5.8) with weights
$(n_{1}\ldots n_{2^{k-1}}m_{1}\cdots m_{2^{k-1}})^{\omega-\frac{1}{k}}(\log n_{1})\cdots(\log n_{2^{k-1}})(\log m_{1})\cdots$ $(\log m_{2^{k-1}})$
$\hskip 30.00005pt\ll\bigg{(}\frac{x^{\omega}}{x^{1/k}/\log x}\bigg{)}^{2^{k}},$
@@
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d\alpha}\ll\frac{x^{2^{k}\omega}}{(x^{1/k}/\log x)^{2^{k}}}\int_{0}^{1}|g(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\ll x^{2^{k}\omega-1}(\log x)^{c(k)+2^{k}}.$
@@
for $\ell\geq 2^{k}$ we get
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{(\ell-2^{k})\omega}\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{c(k)+2^{k}}.\qed$
@@
Since $h\geq 2^{k}+1$, it follows from Lemma 5.13 that
$\int_{\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h}\,\mathrm{d}\alpha$ $\leq\bigg{(}\sup_{\alpha\in\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h-2^{k}}\bigg{)}\,\int_{0}^{1}|\widetilde{T}(\alpha;N)|^{2^{k}}\,\mathrm{d}\alpha$
$\ll\bigg{(}\frac{N^{\omega}}{(\log N)^{4^{1-k}C-G-1}}\bigg{)}^{h-2^{k}}N^{2^{k}\omega-1}(\log N)^{c(k)+2^{k}}$
@@
$\bigg{|}\sum_{j=1}^{h}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,$ $e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}$
@@
$\leq\sum_{j=1}^{h}\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{r(C)}}\bigg{)}^{j/h}\,\big{(}(\log N)^{h+c(k)+2^{k}}N^{h\omega-1}\big{)}^{1-j/h}$
$\ll\frac{N^{h\omega-1}}{(\log N)^{r(C)/h-(h+c(k)+2^{k})(h-1)/h}}.$
Writing $R=R(C):=r(C)/h-(h+c(k)+2^{k})(h-1)/h$, we see that choosing $C$ large, $R$ can be taken to be as large as desired, concluding the proof. ∎
@@ 6. Waring–Goldbach subbases
We are now in shape to prove Theorem 1.3. By the prime number theorem, $|\mathbb{P}^{k}\cap[1,x]|\sim kx^{1/k}/(\log x)$, which is a regularly varying function, so $\mathbb{P}^{k}$ satisfies condition (i) of Theorem 1.4. By Lemma 5.10, $\mathbb{P}^{k}$ also satisfies condition (ii), with $H_{\textnormal{(ii)}}\leq 2^{k-1}$.
For condition (iii), we claim that $H_{\textnormal{(iii)}}\leq 2^{k}+1$ for the functions considered in Theorem 1.3. Let $h\geq 2^{k}+1$, and $F(x)=x^{\kappa}\psi(x)$ be a regularly varying function with
$F(x)\leq(1+o(1))\frac{\Gamma(1/k)^{h}}{\Gamma(h/k)}\frac{x^{h/k-1}}{(\log x)^{h}}$
such that $F(x)(\log x)^{h}/x^{h/k-1}$ asymptotic to a non-increasing function. Note that the upper bound for $F$ comes from the total number of representations of $n$ as a sum of $h$ $k$-th powers of primes (cf. Hua [8, Theorem 11]). Define $f(x):=(xF(x))^{1/h}$, so that $f(x)\ll(1+o(1))\pi(x^{1/k})\asymp x^{1/k}/(\log x)$. For the $\eta=\eta(f,k,h)$ of Proposition 2.2, we have
@@
Choose $F(x):=x^{\kappa}\psi(x)$ for some $10\leq\kappa\leq h/k-$ 1 and some measurable positive real function $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$. Suppose that $\log n\ll F(n)\ll x^{h/k-1}/(\log n)^{h}$. We have $f(x)=x^{(1+\kappa)/h}\psi(x)^{1/h}$, where $1/h\leq(1+\kappa)/h\leq 1/k$. By uniform convergence,555[1, Theorem 2.0.8] applied to $g(x):=\psi(e^{x})$, so that $g(\lambda x)\asymp_{\lambda}g(x)$. it follows that $\psi(x_{i})\asymp\psi(n)$ for $n^{\eta/2}\leq x_{i}\leq n$, so Theorem 1.2 implies that the sum in (6.1) is $\asymp n^{\kappa}\psi(n)$ for $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$. This fits into condition (iii) (with “$\mathfrak{S}_{B,h}(n)$” different from $\mathfrak{S}^{*}_{k,h}(n)$, but still $\asymp 1$).
Since $\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}\leq 2^{k}+1$, the conclusion of Theorem 1.3 follows.

### v2

@@ 5.2.2. Integral over minor arcs
Let $h_{0}$ be as in Lemma 5.10. For $\ell\geq 2h_{0}$, we have
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{\xi+2h_{0}}.$
@@
The integral $\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2\ell}\,\mathrm{d}\alpha$ counts the number of solutions to the Diophantine equation (5.8) with weights
$(n_{1}\ldots n_{\ell}m_{1}\cdots m_{\ell})^{\omega-\frac{1}{k}}(\log n_{1})\cdots(\log n_{\ell})(\log m_{1})\cdots$ $(\log m_{\ell})$
$\hskip 30.00005pt\ll\bigg{(}\frac{x^{\omega}}{x^{1/k}/\log x}\bigg{)}^{2\ell},$
@@
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2\ell}\,\mathrm{d\alpha}\ll\frac{x^{2\ell\omega}}{(x^{1/k}/\log x)^{2\ell}}\int_{0}^{1}|g(\alpha;x)|^{2\ell}\,\mathrm{d}\alpha\ll x^{2\ell\omega-1}(\log x)^{\xi+2\ell}.$
@@
for $\ell\geq 2h_{0}$ we get
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{(\ell-2h_{0})\omega}\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2h_{0}}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{\xi+2h_{0}}.\qed$
@@
Since $h\geq h_{k}^{*}\geq 2h_{0}+1$, where $h_{0}$ is as in Lemma 5.10, it follows from Lemma 5.13 that
$\int_{\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h}\,\mathrm{d}\alpha$ $\leq\bigg{(}\sup_{\alpha\in\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h-2h_{0}}\bigg{)}\,\int_{0}^{1}|\widetilde{T}(\alpha;N)|^{2h_{0}}\,\mathrm{d}\alpha$
$\ll\bigg{(}\frac{N^{\omega}}{(\log N)^{4^{1-k}C-G-1}}\bigg{)}^{h-2h_{0}}N^{2h_{0}\omega-1}(\log N)^{\xi+2h_{0}}$
@@
$\sum_{j=1}^{h}\bigg{|}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,$ $e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}$
@@
$\leq\sum_{j=1}^{h}\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{r(C)}}\bigg{)}^{j/h}\,\big{(}(\log N)^{h+\xi+2h_{0}}N^{h\omega-1}\big{)}^{1-j/h}$
$\ll\frac{N^{h\omega-1}}{(\log N)^{r(C)/h-(h+\xi+2h_{0})(h-1)/h}}.$
Writing $R=R(C):=r(C)/h-(h+\xi+2h_{0})(h-1)/h$, we see that choosing $C$ large, $R$ can be taken to be as large as desired, concluding the proof. ∎
@@ 6. Waring–Goldbach subbases
We are now in shape to prove Theorem 1.3. By the prime number theorem, $|\mathbb{P}^{k}\cap[1,x]|\sim kx^{1/k}/(\log x)$, which is a regularly varying function, so $\mathbb{P}^{k}$ satisfies condition (i) of Theorem 1.4. By Lemma 5.10, $\mathbb{P}^{k}$ also satisfies condition (ii), with $H_{\textnormal{(ii)}}\leq\frac{1}{2}h_{k}^{*}-1$.
For condition (iii), we claim that $H_{\textnormal{(iii)}}\leq h^{*}_{k}$ for the functions considered in Theorem 1.3. Let $h\geq h_{k}^{*}$, and $F(x)=x^{\kappa}\psi(x)$ be a regularly varying function with
$F(x)\leq(1+o(1))\frac{\Gamma(1/k)^{h}}{\Gamma(h/k)}\frac{x^{h/k-1}}{(\log x)^{h}}.$
Note that the upper bound for $F$ comes from the total number of representations of $n$ as a sum of $h$ $k$-th powers of primes (cf. Hua [8, Theorem 11]). Define $f(x):=(xF(x))^{1/h}$, so that $f(x)\ll(1+o(1))\pi(x^{1/k})\asymp x^{1/k}/(\log x)$. For the $\eta=\eta(f,k,h)$ of Proposition 2.2, we have
@@
Choose $F(x):=x^{\kappa}\psi(x)$ for some $10\leq\kappa\leq h/k-$ 1 and some measurable positive real function $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$. Suppose that $\log n\ll F(n)\ll x^{h/k-1}/(\log n)^{h}$. We have $f(x)=x^{(1+\kappa)/h}\psi(x)^{1/h}$, where $1/h\leq(1+\kappa)/h\leq 1/k$. By uniform convergence,333[1, Theorem 2.0.8] applied to $g(x):=\psi(e^{x})$, so that $g(\lambda x)\asymp_{\lambda}g(x)$. it follows that $\psi(x_{i})\asymp\psi(n)$ for $n^{\eta/2}\leq x_{i}\leq n$, so Theorem 1.2 implies that the sum in (6.1) is $\asymp n^{\kappa}\psi(n)$ for $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$. Using the fact that $\psi\asymp\phi$ for some $\phi$ of regular variation (see Remark 6.1), we can apply Theorem 1.4 to $x^{\kappa}\phi(x)$ instead, so that this fits into condition (iii) (with “$\mathfrak{S}_{B,h}(n)$” different from $\mathfrak{S}^{*}_{k,h}(n)$, but still $\asymp 1$).
Since $\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}\leq h^{*}_{k}$, the conclusion of Theorem 1.3 follows.
Remark 6.1.
In order to apply Theorem 1.4 to $F(x)=x^{\kappa}\psi(x)$ with $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$, we need the fact that $\psi(x)\asymp\phi(x)$ for some $\phi(x)$ of slow variation. We are assuming that $\psi(e^{x})\asymp_{\lambda}\psi(e^{\lambda x})$, that is, that $\psi(e^{x})$ is of $O$-regular variation. By the representation theorem for $O$-regular varying functions [1, Theorem 2.2.7], there exist measurable bounded functions $\alpha$, $\xi:\mathbb{R}_{>0}\to\mathbb{R}$ such that
$\psi(e^{x})=\exp\bigg{(}\alpha(x)+\int_{1}^{x}\frac{\xi(t)}{t}\,\mathrm{d}t\bigg{)}.$
Writing $y=e^{x}$, we have
$\psi(y)$ $\asymp\exp\bigg{(}\int_{1}^{\log y}\frac{\xi(t)}{t}\,\mathrm{d}t\bigg{)}=\exp\bigg{(}\int_{1}^{y}\frac{\xi(\log u)}{u\log u}\,\mathrm{d}u\bigg{)}$
Writing $\epsilon(u):=\xi(\log u)/\log u$, since $\xi$ is bounded, $\epsilon(u)\to 0$ as $u\to\infty$. Therefore $\psi(x)\asymp\phi(x):=\exp(\int_{1}^{x}\epsilon(t)/t\,\mathrm{d}t)$, which is a slowly varying function (cf. [1, Theorem 1.3.1]).